Three cards are randomly selected from a deck of 52 cards. After every draw, the card is NOT replaced back in the deck. Find the probability of drawing two Kings in a row, then an Ace.
Answer by greenestamps(13200)



Use the fundamental counting principle.

P(king first) = 4/52
P(king second) = 3/51
P(ace third) = 4/50

Multiply those probabilities to find the probability of getting that sequence of outcomes.